OpenStudy (anonymous):

Need help with Polynomial division and Rewriting expressions!! If x=4y+1 and z=3x-5 find an expression for z in terms of y If (2x^2+kx+3)/(x-3)=2x-1 find the value of k

OpenStudy (anonymous):

You mean substitute it in?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (cwrw238):

substitute x = 4y + 1 in z = 3x - 5 that is replace x in the formula for z by 4y + 1

OpenStudy (cwrw238):

then simplify

OpenStudy (anonymous):

How does it simplify to 12y-2 though?

OpenStudy (cwrw238):

z = 3(4y + 1) - 5 = 12y + 3 - 5 = 12y - 2

OpenStudy (anonymous):

and you did 3*1 which is why it remained "+3" correct?

OpenStudy (cwrw238):

for the second question expand (x - 3)(2x - 1) then compare term by term with 2x^2 + kx + 3

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

by expand you just mean multiply right?

OpenStudy (cwrw238):

distribute 3 over the 4y and the 1 in the parentheses

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

compare term by term?

OpenStudy (anonymous):

Also thank you for the help

OpenStudy (cwrw238):

first multiply second parentheses by x then by -3 then simplify

OpenStudy (cwrw238):

well you'll obtain an expression with 3 terms one in x^2 , one in x and one a number line this up with 2x^2 + kx + 3 and you'll see what k is

OpenStudy (anonymous):

Oh I simplify the number I get from (x - 3)(2x - 1)? than compare it

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

yes